The average hourly wage of employees in a large nationwide industry is $12.96. Assume the variable is normally distributed, with a standard deviation of $4.46. In a randomly selected sample of 10,000 such employees, how many of them (correct to the nearest hundred) would you expect to be earning between $9.00 and $17.00 per hour?
Answer by stanbon(57377)

The average hourly wage of employees in a large nationwide industry is $12.96. Assume the variable is normally distributed, with a standard deviation of $4.46. In a randomly selected sample of 10,000 such employees, how many of them (correct to the nearest hundred) would you expect to be earning between $9.00 and $17.00 per hour?
-----
z(9) = (9-12.96)/4.46 = -0.8879
----
z(17) = (17-12.96)/4.46 = 0.9058
---------------------------------------
P(9 < x < 17) = P(-0.8879 < z < 0.9058) = 0.6302
================
Cheers,
Stan H.
================